If the diameter of the field of view of a compound microscope at 40X magnification is 6,000 micrometers (um), what would be the field of view at 400X magnification?
Answer #1

Diameter of the field = 6000 um

Known magnification = 40X

Unknown magnification = 400X

The field size is inversely proportional to the magnification

The relationship between the total magnification and the diameter of the field of view is inverse. so, as magnifications increases the diameter of the field decreases in proportion, so the diameter of field of view at different magnification can be estimated by using the formula

\frac{Known\ magnification}{Unknown\ magnification}\times original \ diameter=Diameter\ of\ \ Field \ of \ View

\frac{40}{400}\times 6000=600um